# Breakdown of sequential tunnel ionization in ultrashort electromagnetic pulses

### v1

@@ Abstract
We consider double ionization of negative bromine ion ${\rm Br}^{-}$ in intense linearly polarized low-frequency laser fields. By solving numerically the two-electron time-dependent Schrödinger equation we demonstrate that while for electromagnetic pulses of a few tens of femtoseconds duration and longer the sequential single-electron approximation perfectly describes the ionization dynamics, for pulses as short as a few femtoseconds this picture breaks down entirely, and the electron-electron interaction suppresses the rate of ionization by approximately one order of magnitude. We also show weak but clear signatures of the collective tunneling effect in the photoelectron density distribution.
@@
The fundamental effect of quantum mechanical tunneling discussed for the first time in the context of $\alpha$-decays [1, 2] is pivotal in the physics of nuclei [3], the physics of atoms in external fields [4, 5], and the physics of condensed matter [6]. Optical tunneling in atoms, molecules, and solids subject to intense infrared laser fields is the key phenomenon of contemporary nonlinear optics. It underlies attosecond science [7, 8], the generation of high-order harmonics (HHG) [9, 10], and attosecond spectroscopy [11, 12, 13]. Tunnel and multiphoton ionization of atoms in the field of intense laser radiation first theoretically addressed by Keldysh [14], was observed initially as multiquantum absorption [15] in the mid-60es, and later – in the form of above-threshold ionization (ATI) [16], a process where the electron absorbs more photons when necessary to escape from the atom. Both ATI and HHG can be efficiently explained within the semi-classical three-step model [17, 18], where optical tunneling serves as the first step.
@@
A wide range of theoretical approaches describing the phenomenon of strong-field ionization and related processes are based on the single-active electron (SAE) approximation. This is commonly substantiated by the fact that, in the regime of optical tunneling, the electron must travel a sufficient distance from the nucleus to appear in that part of space where the laser force dominates over the Coulomb attraction. The higher the ionization potential $I_{p}$ of the level, the smaller is the probability that the electron appears at large distances from the nucleus, so that the model in which the currently outermost electron solely interacts with the laser field sounds adequate. We will refer to the regime of ionization dominated by the single-electron dynamics as sequential ionization. The celebrated exclusion from this rule takes place in linearly polarized fields of non-relativistic intensity, where the outermost electron may undergo a recollision with the atomic residual along the same scenario, as underlies the effects of HHG [18] and high-energy ATI plateau [19]. Recollision leads to an extreme enhancement in double ionization that is up to several orders of magnitude, compared to that expected on the basis of the sequential mechanism. This effect of nonsequential double ionization (NSDI) is the eminent example of an ionization channel strongly mediated by electron-electron interaction in the regime of tunneling. By now, NSDI in atoms has been thoroughly studied both experimentally and in theory so that its recollision genesis is beyond doubt; see Ref.[20] for a review. Other nosequential ionization mechanisms including the shake-off model [21] and the collective tunneling model [22, 23, 24] have been theoretically discussed and numerically examined [25, 23] but not yet observed in experiments.
@@
In linearly polarized fields of nonrelativistic intensity, the NSDI through recollision dominates other possible ionization channels by orders of magnitude, which makes a search of alternative nonsequential mechanisms extremely difficult both in numerical calculations and in experiment. However, there are three ionization regimes where such alternatives can be revealed: (i) relativistic ionization, (ii) ionization in fields with circular or elliptical polarization where rescattering is damped out be the two-dimensional motion of the laser-driven electron (see e.g. [26, 27] for details and examples), and (iii) ionization by extremely short laser pulses with duration insufficient to bring the electron back to its parent ion. The latter is hardly achievable in experiment, but easy to reach in theoretical consideration, where modeling of strong-field ionization by short unipolar pulses allows for addressing tiny effects otherwise shielded by the oscillating electron’s dynamics [28].
@@
Here, some elaboration of the terminology is required to avoid ambiguities and misinterpretations in the what follows. We characterize ionization as sequential if the electrons escape from the atom one after another with some significant time interval. The electron-electron interaction may or may not play an important role in this process, so that the sequential ionization is not necessarily equivalent to ionization of model non-interacting electrons. The nonsequential ionization is apparently something opposite, so it is a process in which both electrons depart from the ion simultaneously or within a small time window. Note that in the recollision induced NSDI the first electron leaves the atom approximately one laser period earlier than the second one. However, this fact does not withdraw this channel from the present definition, because eventually the electrons depart simultaneously, after the recollision, while the preceding evolution of the first electron is a virtual process. Note that the term NSDI is equally applicable to different mechanisms of double ionization [29, 30, 23, 25]. Simultaneous emission can take place even in the case of non-interacting electrons or for interacting electrons, but without recollision and excitation of intermediate bound states. We will refer to such simultaneous emission as collective tunneling [22, 23].
@@
In this paper, we examine the ionization dynamics of a two-electron atomic system subject to a strong time-dependent electric field under conditions that eliminate recollision. We choose the bromine negative ion as a target with two very different ionization potentials to provide conditions, where the collective channel of ionization [22, 23] can also potentially make a sizable contribution [24]. The Keldysh parameter $\gamma=\sqrt{2mI_{p}}\omega/eE_{0}$ [14] is kept about unity or smaller, so that the ionization proceeds through the tunneling mechanism. Here $\omega$ and $E_{0}$ are the characteristic frequency and strength of the electric field, $I_{p}$ is the ionization potential, $e$ and $m$ are the elementary charge and mass correspondingly. Solving numerically the time-dependent Schrödinger equation (TDSE), we calculate the probability of double ionization for the 2-dimensional (2D) ${\rm Br}^{-}$ ion and show that, very surprisingly, in short pulses with duration $\approx 2$ fs this probability deviates dramatically from that expected in the case of a sequential process, when the electrons are being removed one after another. Not less surprisingly, in longer pulses of several dozens fs duration, a rudimentary sequential tunneling model reproduces the exact numerical result with accuracy close to one per cent. Thus in a single calculation we prove the validity of the sequential tunnel ionization picture for one set of parameters and demonstrate its complete failure for another. Not less importantly, we also show that the two-particle density distribution demonstrates a signature of the collective tunneling effect.
@@
Our calculations made for a 2D two-electron atom, thus effectively for a 4D system, appear demanding as far as the required computer time and memory are concerned. For better justification of our findings we also present some results obtained for a 1D two-electron atom, which allow for tracing the wave function evolution at a much longer time scale.
@@
We compare the ionization dynamics of two model systems: (a) a 2D two-electron negative ion with interacting electrons and (b) a 2D ion with two non-interacting electrons having the same ionization potentials as (a). In case (a), interaction with the parent ion $V_{\text{ion}}$ and electron-electron interaction $V_{\text{int}}$ are described by the smoothed Coulomb potentials (here and below we use atomic units $m=e=\hbar=1$)
@@
The smoothing parameter $a$ is chosen to reproduce the second ionization potential $I_{p2}$ when the outermost electron is removed. After this, the parameter $b$ is adjusted to reproduce the $I_{p1}$ of the outer electron. Parameters $a$ and $b$ and the ground-state wavefunction are found by numerically solving the TDSE (see details below).
The two-electron TDSE for the case (a) of the interacting electrons has the form
where $H_{0}$ is the single particle Hamiltonian including interaction with the electric field of the laser pulse
with $\varphi$ being its scalar potential.
To suppress recollisions, the electric field is set in the form of an unipolar pulse polarized along the $x$-axis with the amplitude value $E_{0}$ and duration $T$
$E(t)=-E_{0}\sin^{2}\left(\dfrac{\pi t}{T}\right)~{}.$ (5)
In this spatially homogeneous field the electrons are accelerated after ionization, rapidly reaching the boundaries of the computation box. This hinders analysis of the ionization dynamics. To make such analysis more affordable, in the numerical calculation we use the field which is homogeneous in the vicinity of the ion and vanishes far from it:

### v2

@@ Abstract
We consider double ionization of negative bromine ion in intense low-frequency electromagnetic fields. By solving numerically the two-electron time-dependent Schrödinger equation we demonstrate that while for pulses of a few tens of femtoseconds duration and longer the sequential single-electron approximation perfectly describes the ionization dynamics, for pulses as short as a few femtoseconds this picture breaks down entirely, and the electron-electron interaction suppresses the rate of ionization by roughly one order of magnitude. We also show clear signatures of the collective tunneling effect in the photoelectron density distribution. This counter-intuitive channel of ionization opens up due to the electron-electron repulsion in the direction lateral to the applied electric field.
@@
The fundamental effect of quantum mechanical tunneling discussed for the first time in the context of $\alpha$-decays [gamow-28, condon-pr29] and ionization of hydrogen in a weak static field [opp_pr28] is pivotal in the physics of nuclei [weisskopf-book], the physics of atoms in external fields [fedorov-book, popov-usp04], and the physics of condensed matter [ankerhold-07]. Optical tunneling in atoms, molecules, and solids underlies attosecond science [ivanov-rmp09, calegari-adv16], the generation of high-order harmonics (HHG) [gaarde-josab22, strelkov-usp23], and attosecond spectroscopy [agostini_nobel24, huillier_nobel24, krausz_nobel24]. Tunnel and multiphoton ionization of atoms in the field of intense laser radiation first theoretically addressed by Keldysh [keldysh-jetp65], was observed initially as multiquantum absorption [delone-jetp66], and later – in the form of above-threshold ionization (ATI) [agostini-prl79]. The effects of ATI and HHG can be efficiently explained within the semi-classical three-step model [kuchiev-jetpl87, corkum-prl93], where optical tunneling serves as the first step.
@@
A wide range of theoretical approaches describing the phenomenon of strong-field ionization and related processes are based on the single-active electron (SAE) approximation. This is commonly substantiated by the fact that, in the regime of optical tunneling, the electron must travel a sufficient distance from the nucleus to appear in that part of space where the laser force dominates over the Coulomb attraction. The higher the ionization potential $I_{p}$ of the level, the smaller is the probability that the electron appears at large distances from the nucleus, so that the model in which the currently outermost electron solely interacts with the laser field sounds adequate. The celebrated exclusion from this rule takes place in linearly polarized fields of non-relativistic intensity, where such an electron may undergo a recollision with the atomic residual along the same scenario, as underlies the effects of HHG [corkum-prl93] and high-energy ATI plateau [paulus-prl94]. Recollision leads to an extreme enhancement in double ionization (DI) that is up to several orders of magnitude, compared to that expected for the sequential mechanism. This effect of nonsequential double ionization (NSDI) is the eminent example of an ionization channel strongly mediated by electron-electron interaction in the regime of tunneling. By now, NSDI in atoms has been thoroughly studied both experimentally and in theory so that its recollision genesis is beyond doubt; see Ref.[becker-rmp12] for a review. Other nosequential ionization mechanisms including the shake-off model [agostini-rev95] and the collective tunneling model [zon-jetp99, becker-prl00, tyurin-leb23] have been theoretically and numerically examined [lein_prl00, becker-prl00] but not yet observed in experiments.
@@
In linearly polarized fields of non-relativistic intensity, the NSDI through recollision dominates other possible DI channels by orders of magnitude, which makes a search for alternative mechanisms extremely difficult both in numerical calculations and in experiment. Moreover, the NSDI associated with recollisions appears a remarkably broad concept, which allows also entirely classical interpretations [eberly_prl05, mauger_prl12, huang_oe13] and covers, in a row with the plane hard recollision followed by the immediate release of the second electron [kopold_prl00], several indirect sub-channels. In particular, the second electron can be emitted through the excitation of a bound state followed by field ionization [rudenko_prl04], through soft recollisions [emma_pra09], or the excitation into continuum can be facilitated by the complex dynamics of the laser-dressed outer electron [mauger_prl12]. There are three DI regimes where alternatives of the recollision scenario can be revealed: (i) relativistic ionization, (ii) ionization in fields with circular or elliptical polarization where rescattering is damped out due to the two-dimensional motion of the laser-driven electron (see e.g. [keitel-pra07, kolya-pra08] for details and examples), and (iii) ionization by extremely short laser pulses with duration insufficient to bring the electron back to its parent ion. NSDI in ultra-short laser pulses has experimentally demonstrated a transition between regimes with few and multiple recollisions [kubel_njp14]. Even shorter pulses supporting no recollision are hardly achievable in experiments, but they are easy to reach by modeling strong-field ionization in unipolar fields. This allows for addressing tiny effects otherwise shielded by the oscillating electron’s dynamics [keitel-pra22].
@@
For the following, some elaboration of the terminology is required to avoid ambiguities and misinterpretations. We characterize ionization as sequential if the electrons escape from the atom one after another with some significant time interval. The electron-electron interaction may or may not play an important role in this process, so that the sequential ionization is not necessarily equivalent to ionization of model non-interacting electrons. The nonsequential ionization is by definition something opposite, so it is a process in which both electrons depart from the ion within a small time window. Note that in the recollision induced NSDI the first electron leaves the atom approximately one laser period earlier than the second one. However, this fact does not withdraw this channel from the present definition, because eventually the electrons depart simultaneously, after the recollision, while the preceding evolution of the first electron is a virtual process. The term NSDI is equally applicable to different mechanisms of DI [kulander_prl92, moshammer_prl01, becker-prl00, lein_prl00]. Simultaneous emission can take place for non-interacting electrons or for interacting ones, but without recollision and excitation of intermediate bound states.
@@
In this paper, we examine the ionization dynamics of a two-electron atomic system subject to a strong time-dependent electric field under conditions entirely eliminating recollision. We choose the bromine negative ion as a target with two very different ionization potentials to provide conditions, where the collective channel of ionization [zon-jetp99, becker-prl00] can also potentially make a sizable contribution [tyurin-leb23]. The Keldysh parameter $\gamma=\sqrt{2mI_{p}}\omega/eE_{0}$ [keldysh-jetp65] is kept about unity or smaller, so that the ionization proceeds through tunneling. Here $\omega$ and $E_{0}$ are the characteristic frequency and strength of the electric field, $I_{p}$ is the ionization potential, $e$ and $m$ are the elementary charge and mass correspondingly. Solving numerically the time-dependent Schrödinger equation (TDSE), we calculate the probability of DI for the 2-dimensional (2D) $\rm Br$$^{-}$ ion and show that in short pulses with duration $\approx 2$ fs this probability deviates dramatically from that expected in the case of a sequential process. Not less remarkably, in longer pulses of several dozens fs duration, a rudimentary sequential tunneling model reproduces the exact numerical result with accuracy of few per cent. Thus in a single calculation we prove the validity of the sequential tunnel ionization picture for one set of parameters and demonstrate its complete failure for another. . Most importantly, we also demonstrate a clear signature of simultaneous emission, which can be persuasively interpreted as collective tunneling.
@@
Our 2D calculations made effectively for a 4D system, appear demanding as far as the required computer time and memory are concerned. For substantiation of our findings, we also present some results obtained for a 1D two-electron atom, which allow for tracing the wave function evolution at a much longer time scale.
@@
We compare the ionization dynamics of two model systems: (a) a 2D two-electron negative ion with interacting electrons and (b) a 2D ion with two non-interacting electrons having the same ionization potentials as (a).
The two-electron TDSE for the case (a) reads (here and below $m=e=\hbar=1$)
where $H_{0}$ is the single particle Hamiltonian including interaction with the electric field of the pulse
with $\varphi$ being its scalar potential. Definitions of the electron-electron and electron-ion interaction potentials $V_{\text{int}}$ and $V_{\text{ion}}$, along with the model of the pulse field are given in the End Matter. The computational grid parameters are provided there, too. In the case (a) the TDSE (1) is solved for the 1D and 2D systems. In the case (b), the TDSE with the single-particle Hamiltonian (2) was solved for the first and second electrons separately (see the End Matter for details). To suppress recollisions, the time dependence of the electric field $E(t)$ is set in the form of an unipolar pulse with the amplitude value $E_{0}$ and the full duration $T$
$E(t)=-E_{0}\sin^{2}\left(\dfrac{\pi t}{T}\right)~.$ (3)
In both cases (a) and (b) the initial wave function $\Psi_{0}(\bm{r}_{1},\bm{r}_{2},t)$ was symmetrized with respect to coordinates assuming the full spin 0. The two-electron ground state wave function was obtained by using the imaginary time method [bader2013]. The double ionization probability $W(t)$ was found by integrating the modulus squared wave function over space with the ion vicinity excluded
$W(t)=\int\limits_{\begin{subarray}{c}|\bm{r}_{1,2}|>\varkappa\end{subarray}}|\Psi(\bm{r}_{1},\bm{r}_{2},t)|^{2}d\bm{r}_{1}d\bm{r}_{2}~.$ (4)
Here, $\varkappa=15$ is the size of the area around the ion, which was selected such that the probability $W$ does not depend on $t$ at the end of the calculation.
$F(x_{1},x_{2},t)=\int\limits_{\begin{subarray}{c}|y_{1}|>5~\text{a.u.}\\ |y_{2}|>5~\text{a.u.}\end{subarray}}|\Psi(x_{1},y_{1},x_{2},y_{2},t)|^{2}dy_{1}dy_{2}~.$ (5)
Cutting out small values of $y_{1}$ and $y_{2}$ eliminates most of the contribution from single ionization (when the second electron is located near $y=0$) and makes the double ionization flux better visible on the background generated by the outer electron.
We consider two ionization regimes. In the first case, we use a long electromagnetic pulse of amplitude $E_{0}=0.035$ (corresponding to the peak intensity $\approx 10^{14}{\rm W/cm}^{2}$), and a duration of T=40T=40fs. Below we refer this case to as the “long pulse regime” (LPR). In the LPR the ionization probability for the inner electron becomes sizable, when the outer orbital is already depleted. In this case, the electron-electron interaction is significant in the bound state only, and we expect the SAE approximation describing ionization with good accuracy. FIG. 1 shows the two-electron probability distribution for the interacting (a) and non-interacting (b) electrons in 1D. These distributions look rather similar, except for a small population near the diagonal $x_{1}=x_{2}$ present for the non-interacting electrons (b). The general similarity of this distributions indicates that, in the case of sufficiently long pulses, e-e interaction does not significantly affect the ionization process, because the most probable instants of detachment of the outer and the inner electrons are well separated in time. To provide a quantitative measure of this similarity, we integrate the populations over space to find the total probability of DI (4). It appears remarkably insensitive to the presence of e-e interaction. Namely, this probability changes by about $2\%$ if the interaction is taken into account. (FIG. 2(a), solid lines). A similar calculation with the equally long pulse for 2D electrons is unaffordable in view of the computer resources, so we reduced the pulse duration to $T=13$ fs. The distribution function calculated for the 2D systems (5) is shown in Fig.1, panels (c) and (d). Qualitatively, they resemble the distributions of (a), (b) with the only significant difference, which is the presence of a tiny diagonal population for the interacting electrons (c). The total probability of DI (4) for the interacting and non-interacting 2D electrons (FIG. 2(b)) differ more than in the 1D case (FIG. 2(a)), but the agreement is pretty good, and we expect that the difference will decrease with the pulse duration increasing.
Figure 1: Probability density $|\Psi(x_{1},x_{2},t)|^{2}$ for the 1D ion (a,b) and distribution $F(x_{1},x_{2},t)$ calculated along (5) for the 2D ion (c,d) with the interacting (a,c) and non-interacting electrons (b,d). The distributions are shown for the center of the pulse ($t=T/2$).
Figure 2: Total probability of double ionization (4) in the LPR (a) and SPR (b). Blue and red lines correspond to the 1D and 2D cases respectively. Solid lines show the case of the interacting electrons, dashed lines – that of the non-interacting electrons. The field parameters: $E_{0}=0.035$, T=40T=40fs and T=13T=13fs (a), T=2T=2fs (b).
Now we turn to the criterion of separation between the LPR and the short pulse regime (SPR). Fig.3 shows the populations $n_{1}$ of the single ionized species (neutral bromine) and $n_{2}$ of the doubly ionized species $\rm Br$$^{+}$ as functions of time. These results were obtained by numerically solving the system of rate equations
[…]
where $n_{0}$ is the population of the seeding specie (negative ion $\text{Br}^{-}$ in our case). For the time-dependent ionization rates $w_{1,2}[E(t)]$ in (6), we use the sequential tunnel Perelomov-Popov-Terent’ev (PPT) formulas [popov_jetp66, perelomov_jetp67] (see reviews [popov-usp04, poprz_jpb14] for detailed descriptions). For the present case, we are not addressing effects of the bound state angular momentum on the rate of ionization, so $l=m=0$ was set in the PPT rates. Fig. 3 illustrates the qualitative difference between the LPR (a) and SPR (b): in the former case the outer orbital is entirely depleted before the inner electron can be considerably affected by the laser field; in the SPR the inner electron begins to feel the field before the outer one is removed. The border between the two regimes is determined by the ionization saturation time $t_{s}$ for the outer electron. Here we determine this time as such pulse duration that the population $n_{1}$ of singly ionized atoms reaches 0.8 at the field maximum. Numerical analysis of (6) gives $t_{s}\approx 12$ fs for $E_{0}=0.035$. This suggest that for the pulse duration $T\approx 2$ fs the SPR is amply achieved (Fig. 3(b)). Both in the LPR and SPR the Keldysh parameter estimated for the effective frequency $\omega=\pi/T$ falls either into the intermediate (for T=2T=2fs for the first ionization potential $\gamma_{1}=0.54$ and for the second one $\gamma_{2}=1.01$) or into the deep tunnel ($\gamma_{1}=0.25$, $\gamma_{2}=0.05$ for T=40T=40fs) regimes, which justifies applicability of the PPT rates. Note however that the semiclassical rates of ionization [perelomov_jetp67] overestimate the probability with the relative intensity increasing [popov-usp04], see the Supplementary Material (SM) for details.
Figure 3: Time-dependent populations of singly (solid black lines) and doubly (dashed black lines) ionized atoms in the LPR (a) and SPR (b).
Figures 4 and 2(b) present the SPR analogues of Figs. 1 and 2(a), respectively. Their pairing comparison reveals both similarities and striking differences in the ionization dynamics. In Fig. 4(b,d) and Fig. 1 (b,d) we see that for the SPR in absence of the e-e interaction a significant part of the electron density is concentrated near the diagonal $x_{1}=x_{2}$, i.e. the electrons tend to escape simultaneously much more than they do in the LPR (as it was illustrated above by solving the rate equations). Fig. 4 (a,c) shows that the e-e interaction suppress this diagonal flux. This leads to a great decrease in DI (see Fig.2(b)): its total probability drops by approximately one order of magnitude when the e-e interaction is taken into account. The suppression of the diagonal flux in the 1D case is almost total (compare Fig. 4(a) and (b)). This can be qualitatively explained by the following naive 1D picture. Let two bound electrons be shifted by the electric field along its direction, and their positions are different. The electron, which is closer to the core, is pulled back not only by the nucleus attraction but also by the repulsion of the other electron. This repulsion reduces the probability of detachment for the inner electron until the outer one is far from the atom.
Figure 4: Same as in Fig.1, but in the short pulse regime. The field parameters and the time instants are: $t=2.5$ fs (a,b), $t=2.0$ fs (c,d), $E_{0}=0.035$, $T=2$ fs.
However, in the 2D case some diagonal flux survives (compare Fig. 4(c) and (d)). To identify the mechanism responsible for this flux, we examine the two-electron motion along the lateral coordinate $y$. Fig.5(a) shows that for the diagonal configuration $x_{1}=x_{2}$, the spatial distribution of the detached electrons exhibits pronounced peaks at $y_{1}=-y_{2}$. This observation, extracted from the ab initio calculation, allows for a clear semiclassical interpretation: the field (3) drives both electrons jointly along the $x$-axis ($x_{1}\approx x_{2}$), while the Coulomb repulsion leads to a transverse separation ($y_{1}\approx-y_{2}$). To determine whether this correlated motion proceeds through tunneling or the over-the-barrier ionization mechanism (in the latter case a simultaneous departure could be interpreted as resulting from laser-driven acceleration), we examine the potential energy profile along trajectories with
$x_{1}=x_{2}=x,~y_{1}=-y_{2}=y~.$ (7)
Figure 5: (a): The normalized spatial distribution $|\Psi(x,y_{1},x,y_{2})|^{2}$ at $x=16.5$, corresponding to the local maximum of the diagonal flux in Fig.4(c) for $t=2.0$ fs; (b): the effective potential barrier (8).
At the maximum of the field (3), the effective potential barrier along the trajectories satisfying (7) takes the form
$U(x,y)=2V_{\text{ion}}(x,y)+V_{\text{int}}(x,y;x,-y)-2E_{0}x+I_{p}~,$ (8)
where $I_{p}=I_{p1}+I_{p2}$ is the total ionization energy, $E_{0}=0.035$. Fig.5(b) demonstrates that all trajectories satisfying (7) encounter a potential barrier separating the area where the initial state is bound and the continuum, confirming the tunneling nature of this ionization channel, see the SM for further details. Crucially, the barrier minimizes its width exactly through the lateral electron separation. This explains why signatures of the collective tunneling are absent in our and earlier [becker-prl00, lein_prl00] 1D TDSE calculations. Note, that the picture of collective tunneling we elaborate here differs fundamentally from that introduced in [zon-jetp99, becker-prl00], where the electrons were assumed to tunnel along a straight 1D line a single quasi-particle with neglect of the Coulomb repulsion.
In conclusion, we numerically examined the DI of a 2D atomic system identifying conditions at which the e-e interaction influences the DI rate, destroying the generally accepted picture of sequential single-electron tunneling. Our results show that this happens in short pulses, when the ionization rate for the inner electron becomes sizable before the outer electron is removed with guarantee. In this case, the DI probability is suppressed by about one order of magnitude, compared to the case of independent sequential tunneling. Thus, the single active electron approximation customary in the theory of strong field processes, is highly limited for intense ultrashort laser pulses.
Most importantly, we demonstrated clear signatures of collective tunneling. We presented a simple picture of simultaneous tunneling in 2D geometry showing that the very same e-e repulsion, which prevents it in 1D [becker-prl00, lein_prl00], generates specific symmetric trajectories along which the electron pair may escape with probability comparable to that of sequential ionization. Thus the target geometry appears pivotal for the very emergence of the effect.
Our predictions may have an important impact at least in two research areas of strong-field atomic physics. Firstly, they suggest the way to a search for collective tunneling in atomic systems subject to tailored laser pulses, which support no recollision. Secondly, the predicted breakdown of the sequential tunneling picture should play a quantitatively important role in the process of multiple ionization of atoms in short laser pulses of extreme intensities, where, due to the high peak field and the steepness of the time envelope the ionization saturation time could lie in the sub-fs domain. In particular, the suppression of ionization induced by the e-e correlation should be pivotal for the development of numerical particle-in-cell codes aimed at modeling the laser-plasma dynamics with multiple field ionization of atoms [derouillat_cpc2018].
End Matter
This End Matter explains details of the TSDE (1, 2). These include the form of the interaction potentials and parameters of the computational grid. In the case of two-electron negative ion with interacting electrons (a) the electron-electron interaction $V_{\text{int}}$ and the interaction with the parent ion $V_{\text{ion}}$ were approximated by the smoothed Coulomb potentials:
@@
The smoothing parameter $a$ is chosen to reproduce the second ionization potential $I_{p2}$ when the outermost electron is removed. After this, the parameter $b$ is adjusted to reproduce the $I_{p1}$ of the outer electron. Parameters $a$ and $b$ and the ground-state wavefunction are found by numerically solving the TDSE. To reproduce the first $I_{p1}=3.37$ eV and second $I_{p2}=11.81$ eV ionization potentials of $\rm{Br}^{-}$ we used $a=1.66$, $b=2.6$ for 1D electrons and $a=1$, $b=2.2$ for 2D electrons in (9) and (10).
In the 1D system, coordinates $x_{1}$ and $x_{2}$ of the electrons were discretized on a spatial grid of $2048\times 2048$ nodes with a step $\Delta x=0.5$ a.u. The absorbing layer [neuhasuer1989] occupies 15 nodes at the boundary of the computational box. For the 2D problem, the coordinates were discretized with a step $\Delta x=\Delta y=0.5$ a.u. and $N_{x}=512$, $N_{y}=256$ nodes for each electron. The total size of the computational box is $512\times 256\times 512\times 256$ nodes.
In the case of an ion with two non-interacting elec- trons (b), the TDSE with the single-particle Hamiltonian (2) was solved for the first and second electrons separately. For the outer electron, the potential was approximated by a short-range well of the form
$V_{\text{short}}(r)=-c_{1}\exp\left(-\dfrac{r^{2}}{c_{2}^{2}}\right)$ (11)
with two fitting parameters $c_{1}$ and $c_{2}$ chosen to reproduce the first ionization potential $I_{p1}$. For the second electron, we used the smoothed Coulomb potential (9) with a fitting parameter $a$ chosen to reproduce $I_{p2}$. The parameters were taken $c_{1}=0.48$, $c_{2}=0.76$ and $a=1.66$ for the 1D problem and $c_{1}=0.48$, $c_{2}=2$ and $a=1$ for the 2D problem.
In Eqs. (1, 2), interaction of the electrons with the field of the unipolar pulse is introduced through the scalar potential: $\varphi(\bm{r},t)=-\int\tilde{E}(x,t)dx$, here $\tilde{E}(x,t)$ is the electric field linearly polarized in the x-direction In a spatially homogeneous field as (3), the electrons are accelerated after ionization, rapidly reaching the boundaries of the computation box. This hinders analysis of the ionization dynamics. To make such analysis more affordable, in the numerical calculation we use the field which is homogeneous in the vicinity of the ion and vanishes far from it: